OpenStudy (skarlettrose186):

Find the average rate of change of the function defined by the following table. x y -2 -5 2 35 6 75 10 115 -1/10 1/10 10 -10

Vocaloid (vocaloid):

let's pick two points on the table (-2, -5) (2, 35) average change = change in y/change in x

Vocaloid (vocaloid):

so, we first subtract the y-coordinates to find the change in y 35 - (-5) = ?

OpenStudy (skarlettrose186):

40?

Vocaloid (vocaloid):

good now we subtract the x-coordinates 2 - (-2) = ?

OpenStudy (skarlettrose186):

4

Vocaloid (vocaloid):

good! so change in y/change in x = 40/4 = 10

OpenStudy (mathmale):

Sorry for having closed your post by mistake. Mea culpa. Just divide the change in y by the change in x to get the average rate of change. This is very similar to finding the slope of a line segment connecting two points.

OpenStudy (skarlettrose186):

ok thank you i was wondering why it got closed

OpenStudy (mathmale):

Not due to any fault of yours.